the sum of two numbers is 42 and their difference is 6 what are the two numbers?
Answer by mananth(16946)

x+y=42 .............1
x-y=6 .............2
Add the two equations

2x=48

/2

x=24

plug value of x in (1)

x+y= 42
24+y=42
y=18